# Smoothing of Headland Path Edges and Headland-to-Mainfield Lane Transitions Based on a Spatial Domain Transformation and Linear Programming

### v1

@@ Abstract
Within the context of in-field path planning and under the assumption of nonholonomic vehicle models this paper addresses two tasks: smoothing of headland path edges and smoothing of headland-to-mainfield lane transitions. Both tasks are solved by a two-step hierarchical algorithm. The first step differs for the two tasks generating either a piecewise-affine or a Dubins reference path. The second step leverages a transformation of vehicle dynamics from the time domain into the spatial domain and linear programming. Benefits such as a hyperparameter-free objective function and spatial constraints useful for area coverage gaps avoidance and precision path planning are discussed. The method, which is a deterministic optimisation-based method, is evaluated on a real-world field solving 3 instances of the first task and 16 instances of the second task.
@@ 1 Introduction
Logistics play a large role in agriculture. For arable farming and the cultivation of crops it can be differentiated between in-field and inter-field logistics (Plessen (2019a)). For the former, efficient path planning is of interest subject to three main considerations: (i) minimisation of tractor traces within the field in order to maximize growable field area, (ii) minimisation of area coverage pathlength in order to minimize fuel consumption and time spent in the field, and (iii) avoiding area coverage gaps such that the entire field is covered, for example during a field-run spraying herbicide or pesticide applications, while simultaneously accounting for a given operating width of the machinery operating in the field and accounting for its typically nonholonomic vehicle dynamics, state and actuator constraints.
@@
For background, this paper is an extension of previous work by the author, in particular, it refines edges of path plans generated in Plessen (2019b) and Plessen (2018) which focused on the higher-level optimisation for the sequence of mainfield lane traversals. See also Höffmann et al. (2024), Mier et al. (2023), Khan et al. (2016) for more high-level area coverage techniques. This paper differs from Plessen & Bemporad (2017), where an online closed-loop tracking controller based on model predictive control was presented, in that here the focus is on offline generation of precision paths, which might then afterwards be tracked online by the approach from aforementioned paper or an alternative closed-loop tracking controller.
@@
The research gap and motivation for this paper is discussed. First, the focus of aforementioned pattern-based references is on lane-to-lane transitions. Second, an issue for pattern-based path planning, in particular based on straights, arcs and clothoid segments, is that resulting paths are typically operating at actuation limits such as maximum curvature and maximum steering rate. For general headland path smoothing (instead of aforementioned lane-to-lane transitions) this is not sufficient where other constraints such as area coverage gaps must also be accounted for. Third, the proposal of a spatial domain transformation in order to be able to formulate optimisation problems with spatial constraints, which is desirable for precision agriculture applications, is missing in the literature.
The focus of this paper is on the path planning step instead of path tracking. The allure of a numerical optimisation approach is that it allows incorporation of a variety of constraints in a disciplined manner and can handle the multivariate nature of the task. The allure of transforming vehicle dynamics from the time into the spatial domain is that it allows to formulate spatial precision constraints. An early paper discussing benefits of transformations of dynamics from the time into the spatial domain is Pfeiffer & Johanni (1987). This paper builds on previous work (Plessen et al. (2017a) and Plessen et al. (2017b)) for autonomous driving. To the author’s knowledge this is the first paper that proposes a spatial transformation of system dynamics to calculate precision path plans in agriculture.
The remaining paper is organised as follows: problem formulation, modeling and proposed solution, numerical results and the conclusion are described in Sections 2-5.
@@ 2 Problem Formulation
Fig. 2: Illustration of terminology and visualisation of Problem 1 and 2.
@@
Given a headland path determine edgy path segments along it and correct them such that a smoothed headland path is generated that is (i) suitable for nonholonomic vehicle dynamics and (ii) avoiding area coverage gaps.
@@
Given headland-to-mainfield lane transitions, and vice versa mainfield lane-to-headland transitions, as part of a field covering path, correct those if necessary such that smoothed transitions are generated that are (i) suitable for nonholonomic vehicle dynamics and (ii) avoiding area coverage gaps.
@@ 3.1 Spatial Domain Transformation
To illustrate the spatial domain transformation and for simplicity, the method is illustrated for a kinematic bicycle model. It is refered to Plessen et al. (2017a) for handling of more complex vehicle models.
This section summarizes the notation of Plessen et al. (2017b). Consider a global and a path-aligned coordinate frame within the $(x,y)$- and $(s,e_{y})$-plane, respectively, with path centerline coordinate $s\geq 0$ and lateral deviation $e_{y}\in\mathbb{R}$. Coordinates can be transformed by projecting point masses (vehicle’s center of gravity) to piecewise-affine path segments. As a detail, it can be differentiated between the path centerline coordinate $s\geq 0$, and actual path length coordinate $\eta\geq 0$, indicating the distance traveled by the vehicle. Note that $\eta\neq s$, unless the vehicle is traveling perfectly along the path centerline. A trajetory can then be defined as $z(s)=\begin{bmatrix}e_{\psi}(s),~{}e_{y}(s)\end{bmatrix}$, with heading deviation $e_{\psi}(s)\in[0,2\pi]$ at coordinate $s$. The equivalent trajectory in the $(x,y)$-plane is defined by $\mathcal{X}(s)=\begin{bmatrix}x(s),~{}y(s),~{}\psi(s)\end{bmatrix}$. The classic nonlinear kinematic bicycle model Rajamani (2011) is
@@
Importantly, (2) is entirely independent of vehicle speed $v$. This is characteristic for kinematic models, but not the case for dynamic vehicle models (Plessen et al. (2017a)). Furthermore, note that (2) has one state dimension less than (1). To summarize, we define spatially-dependent state and control vectors as $z=\begin{bmatrix}e_{\psi},~{}e_{y}\end{bmatrix}$ and $u=\begin{bmatrix}\delta\end{bmatrix}$, respectively. We abbreviate (2) by $z^{\prime}=f(z,u)$. Let a discretisation grid along the path centerline be defined by $\{s_{j}\}_{j=0}^{N}=\{s_{0},s_{1},\dots,s_{N}\}$. For the remainder of this paper, index $j$ refers to the spatial discretisation grid. The discretisation grid can be uniformly spaced. Additional grid points can be added. For a given set of references $\{e_{\psi,j}^{\text{ref}}\}_{j=0}^{N}$, $\{e_{y,j}^{\text{ref}}\}_{j=0}^{N}$ and $\{u_{j}^{\text{ref}}\}_{j=0}^{N-1}$ and initial state, the linearised and discretised (zero-order hold) state dynamics of (2) can be computed for $\{e_{\psi,j}\}_{j=0}^{N}$ and $\{e_{y,j}\}_{j=0}^{N}$ as a function of control inputs $\{\delta_{j}\}_{j=0}^{N-1}$.
@@
whereby the bounds, $\dot{\delta}^{\text{min}}$ and $\dot{\delta}^{\text{max}}$, in general, are nonlinear functions of the vehicle’s operating point and time-varying parameters. By applying the spatial coordinate transformation, a discretisation, a change of variables, and assuming the bounds to remain constant over the planning horizon, we obtain
@@
whereby we abbreviated $D_{s,j}=s_{j+1}-s_{j}$. Thus, linear rate constraints in (3) are rendered not only nonlinear, but also state-dependent and also velocity-dependent. This has two implications. First, to formulate linearly constrained optimisation problems, in general, we require the linearisation of (4). Dependent on the quality of underlying reference trajectories, this may incur significant distortions. Second, while the discrete form of (3) can always be guaranteed to be feasible (assuming a feasible initialisation), for (4) this is not the case anymore. Let us consider two dregrees of simplification of (4). First, we assume small $e_{\psi,j}$ and $e_{y,j}$, and consequently approximate $\frac{(\rho_{s,j}-e_{y,j})}{\rho_{s,j}\cos(e_{\psi,j})}\approx 1$, thereby rendering (4) state-independent, but maintaining velocity-dependent bounds. Thus, steering rate constraints still depend on $v_{j}$. Second, we additionally eliminate this velocity-dependency and formulate
@@
whereby, $v_{\text{ref}}>0$ is introduced as a reference velocity parameter choice. Formulation (5) bears the advantage of eliminating state dependency and is therefore our preferred form for spatial rate constraints. In practice, we employ bounds that are constant over the spatial planning horizon, but in general time-varying in any closed-loop receding horizon control-setting and dependent on the vehicle’s operating point.
To summarize this section, the transformation of time-dependent control rate constraints (3) to the path-aligned coordinate frame is not trivial and to be considered as the main disadvantage of a spatial-based system representation. We opted for the simple state-independent form (5) for linearly constrained optimisation problems, and discussed the role of $v_{\text{ref}}>0$ as a transformation parameter.
@@ 3.2 Main Two-Step Hierarchical Algorithm
An edgy path segment can either be a headland path segment or a headland-to-mainfield lane transition. Edgy path segments are identified by heuristics, in particular, by searching for abrupt heading-changes above a certain threshold as well as the labeling of path coordinates as headland path or mainfield lane and detecting their changes.
@@ 3.3 Problem 1: Headland path smoothing
The two steps of the hierarchical algorithm for Problem 1 are visualised in Fig. 3. The illustrative edgy headland path segment is the one labeled with Problem Type 1 in Fig. 2. First, this segment is scaled by heuristics and replaced by a 5-point piecewise-affine (PWA) path as shown in Fig. 3(a). Underlying notion for this scaling is that the tip of the resulting 5-point PWA path ensures zero spraying gap. Thus, the distance between tip and its projection point along the field contour must be less or equal than half the machinery operating width $w>0$.
As a detail, we do not use popular Chaikin’s algorithm (see Chaikin (1974)) for corner cutting, since this would replace every edge point with two edge points at each iteration step. Furthermore, it would cut off a significant portion of the tip of the 5-point PWA line after the first iteration.
Using the 5-point PWA path from Step 1 as reference path for Step 2 of the hierarchical algorithm and Problem 1 the following basic LP is proposed:
@@
In practice, constraints (6b) which enforce the resulting path to stay on one side of the reference path are relaxed to, $e_{y,j}-e_{y,j}^{\text{ref}}\leq\sigma$, by introducing a non-negative slack variable $\sigma\geq 0$ and adding it to the objective function 6a with a high weight such that $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|$ is replaced with $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|+10^{16}\sigma$. Furthermore, for the absolute value arguments $N$ non-negative surrogate variables are introduced and the set of inequality constraints is extended accordingly. In the end, there are $n_{u}=2N+1$ scalar real-valued optimisation variables. Typical sizes of the LP (i.e., the number of variables and inequality constraints) are stated in the numerical example Section 4.
@@
Finally, LP-iterations or refinement steps can be conducted. Thus, for the first solution of (6) a 5-point-PWA path is used as reference. Then, the resulting path is used as reference for the second solution of (6). In theory, many of such LP-iterations can be carried out. In practice, one refinment step was found to be useful and sufficient.
@@ 3.4 Problem 2: Headland-to-mainfield lanes transitions
Fig. 5 is of special interest for two reasons. First, it illustrates how Dubins paths are not suited for the generation of precision paths to be followed by autonomous machines online, especially when calculating Dubins paths with minimal turning radius of the machine (see Fig. 5(a)). Due to actuation constraints deviations would be expected in any online tracking application. Second, it illustrates the effect on tracking performance when using different turning radii $R_{\text{Dubins}}$ in the Dubins reference path generation step. When using the minimal turning radius Fig. 5(a) results. By using an artificially larger $R_{\text{Dubins}}$ the LP-solution trajectory follows more closely the reference path, see Fig. 5(b) and 5(c). The drawback of a larger $R_{\text{Dubins}}$ is that the resulting path is longer and more compacted area occurs from corresponding tyre traces and the corresponding steering maneuvers, which reduces growable crop area in the field area. A trade-off with a preference towards small $R_{\text{Dubins}}$ for the reference path generation step is recommended.
@@ 4 Numerical Results and Extending Discussion
Fig. 7: Experiment: 3 instances of Problem 1 and 16 instances of Problem 2 needed to be solved. The path plan obtained from proposed method is visualised. The inter-lane distance is 20m.
@@
For numerical evaluation the field from Fig. 1 is considered. It has a size of 4.5ha and requires the solution of 3 instances of Problem 1 and 16 instances of Problem 2. For the purpose of review and evaluation, it is stressed that proposed method is deterministic. The field under consideration involves freeform mainfield lanes, and multiple tight edges for transitions to and from the mainfield lanes and along the headland path. For these reasons and in order to compactly stress characteristics (transition shapes, solve times, lateral deviations) visually and in detail, illustration of results are on a single field for this paper.
@@
Fig. 8 shows a zoom-in for an alternative solution approach to Problem 1. Here, a Bezier-curve of third order is fitted (Pastva (1998)). This requires the solution of a least-squares problem to identify parameters before the evaluation of a Bernstein polynomial. While the interpolating fit is smooth this method does not take vehicle actuation limits into account. The resulting path is therefore not feasible to drive.
@@
4.1 Extending Discussion
By counter-examples and discussion, this subsection is meant to provide further supporting arguments for the suitability of proposed method.

### v3

@@ Abstract
Within the context of in-field path planning and under the assumption of nonholonomic vehicle models this paper addresses two tasks: smoothing of headland path edges and smoothing of headland-to-mainfield lane transitions. Both tasks are solved by a two-step hierarchical algorithm. The first step differs for the two tasks generating either a piecewise-affine or a Dubins reference path. The second step leverages a transformation of vehicle dynamics from the time domain into the spatial domain and linear programming. Benefits such as a hyperparameter-free objective function and spatial constraints useful for area coverage gaps avoidance and precision path planning are discussed. The method, which is a deterministic optimisation-based method, is evaluated on 5 real-world fields solving 19 instances of the first task and 84 instances of the second task.
@@ 1 Introduction
Logistics play a large role in agriculture. For arable farming and the cultivation of crops it can be differentiated between in-field and inter-field logistics (Plessen (2019a)). For the former, efficient path planning is of interest subject to three main considerations: (i) minimisation of tractor traces within the field in order to maximise growable field area, (ii) minimisation of area coverage pathlength in order to minimise fuel consumption and time spent in the field, and (iii) avoiding area coverage gaps (unsprayed areas) such that the entire field is covered, for example during a field-run spraying herbicide or pesticide applications, while simultaneously accounting for a given operating width of the machinery operating in the field and accounting for its typically nonholonomic vehicle dynamics, state and actuator constraints.
@@
Thus, given an optimised sequence of mainfield lane traversals for field coverage such as produced in Plessen (2019b) and Plessen (2018), this paper provides a method for path smoothing. See also Höffmann et al. (2024), Mier et al. (2023), Khan et al. (2016) for more high-level area coverage techniques. This paper differs from Plessen & Bemporad (2017), where an online closed-loop tracking controller based on model predictive control was presented, in that here the focus is on offline generation of precision paths, which might then afterwards be tracked online by the approach from aforementioned paper or an alternative closed-loop tracking controller.
@@
The research gap and motivation for this paper is discussed. First, the focus of aforementioned pattern-based references is on lane-to-lane transitions. Second, an issue for pattern-based path planning, in particular based on straights, arcs and clothoid segments, is that resulting paths are typically operating at actuation limits such as maximum curvature and maximum steering rate. For general headland path smoothing (in contrast to aforementioned lane-to-lane transitions) this is not sufficient where other constraints such as area coverage gaps must also be accounted for. Third, the proposal of a spatial domain transformation in order to be able to formulate optimisation problems with spatial constraints, which is desirable for precision agriculture applications, is missing in the literature.
The focus of this paper is on path planning instead of path tracking. The allure of a numerical optimisation approach is that it allows incorporation of a variety of constraints in a disciplined manner and can handle the multivariate nature of the task. The allure of transforming vehicle dynamics from the time into the spatial domain is that it allows to formulate spatial precision constraints. An early paper discussing benefits of transformations of dynamics from the time into the spatial domain is Pfeiffer & Johanni (1987). To the author’s knowledge this is the first paper that proposes a spatial transformation of system dynamics to calculate precision path plans in agriculture.
The remaining paper is organised as follows: problem formulation, modeling and proposed solution, numerical results and the conclusion are described in Sections 2-6.
@@ 2 Problem Formulation
Fig. 2: Illustration of terminology and visualisation of Problem Type 1 and 2.
@@
Given a headland path, identify path segments with sharp corners or abrupt directional changes, and modify these segments to produce a smoothed headland path that (i) is compatible with nonholonomic vehicle dynamics and (ii) avoids any area coverage gaps.
@@
Given headland-to-mainfield lane transitions, and vice versa mainfield lane-to-headland transitions, as part of a field covering path, correct those if necessary such that smoothed transitions are generated that (i) are compatible with nonholonomic vehicle dynamics and (ii) avoid any area coverage gaps.
@@ 3 Proposed Solution
[…]
Fig. 3: The kinematic bicycle model including the representation of the path-aligned coordinate system.
@@ 3.1 Spatial Domain Transformation
To illustrate the spatial domain transformation and for simplicity, the method is illustrated for a kinematic bicycle model as shown in Fig. 3. It is referred to Plessen et al. (2017a) for handling of more complex vehicle models.
This section summarises the notation of Plessen et al. (2017b). Consider a global and a path-aligned coordinate frame within the $(x,y)$- and $(s,e_{y})$-plane, respectively, with path centerline coordinate $s\geq 0$ and lateral deviation $e_{y}\in\mathbb{R}$. Coordinates can be transformed by projecting point masses (vehicle’s center of gravity) to piecewise-affine path segments. As a detail, it can be differentiated between the path centerline coordinate $s\geq 0$, and actual path length coordinate $\eta\geq 0$, indicating the distance traveled by the vehicle. Note that $\eta\neq s$, unless the vehicle is traveling perfectly along the path centerline. A trajetory can then be defined as $z(s)=\begin{bmatrix}e_{\psi}(s),~{}e_{y}(s)\end{bmatrix}$, with heading deviation $e_{\psi}(s)\in[0,2\pi]$ at coordinate $s$. The equivalent trajectory in the $(x,y)$-plane is defined by $\mathcal{X}(s)=\begin{bmatrix}x(s),~{}y(s),~{}\psi(s)\end{bmatrix}$. The classic nonlinear kinematic bicycle model Rajamani (2011) is
@@
Importantly, (2) is entirely independent of vehicle speed $v$. This is characteristic for kinematic models, but not the case for dynamic vehicle models (Plessen et al. (2017a)). Furthermore, note that (2) has one state dimension less than (1). To summarise, we define spatially-dependent state and control vectors as $z=\begin{bmatrix}e_{\psi},~{}e_{y}\end{bmatrix}$ and $u=\begin{bmatrix}\delta\end{bmatrix}$, respectively. We abbreviate (2) by $z^{\prime}=f(z,u)$. Let a discretisation grid along the path centerline be defined by $\{s_{j}\}_{j=0}^{N}=\{s_{0},s_{1},\dots,s_{N}\}$. For the remainder of this paper, index $j$ refers to the spatial discretisation grid. The discretisation grid can be uniformly spaced. Additional grid points can be added. For a given set of references $\{e_{\psi,j}^{\text{ref}}\}_{j=0}^{N}$, $\{e_{y,j}^{\text{ref}}\}_{j=0}^{N}$ and $\{u_{j}^{\text{ref}}\}_{j=0}^{N-1}$ and initial state, the linearised and discretised (zero-order hold) state dynamics of (2) can be computed for $\{e_{\psi,j}\}_{j=0}^{N}$ and $\{e_{y,j}\}_{j=0}^{N}$ as a function of control inputs $\{\delta_{j}\}_{j=0}^{N-1}$.
@@
whereby the bounds, $\dot{\delta}^{\text{min}}$ and $\dot{\delta}^{\text{max}}$, in general, are nonlinear functions of the vehicle’s operating point and time-varying parameters. By applying the spatial coordinate transformation $\delta^{\prime}=\frac{\dot{\delta}}{\dot{s}}$, a discretisation, a change of variables, and assuming the bounds to remain constant over the planning horizon, we obtain
@@
whereby we abbreviated $D_{s,j}=s_{j+1}-s_{j}$. Thus, linear rate constraints in (3) are rendered not only nonlinear, but also state-dependent and also velocity-dependent. This has two implications. First, to formulate linearly constrained optimisation problems, in general, we require the linearisation of (4). Dependent on the quality of underlying reference trajectories, this may incur significant distortions. Second, while the discrete form of (3) can always be guaranteed to be feasible (assuming a feasible initialisation), for (4) this is not the case anymore. Let us consider two dregrees of simplification of (4). First, we assume given reference trajectories that permit small errors $e_{\psi,j}$ and $e_{y,j}$, and consequently approximate $\frac{(\rho_{s,j}-e_{y,j})}{\rho_{s,j}\cos(e_{\psi,j})}\approx 1$, thereby rendering (4) state-independent, but maintaining velocity-dependent bounds. Thus, steering rate constraints still depend on $v_{j}$. Second, we additionally eliminate this velocity-dependency and formulate
@@
whereby, $v_{\text{ref}}>0$ is introduced as a reference velocity parameter choice. Note that approximation (5) is the exact equivalent of (4) for the zero-reference error setting $e_{y,j}=0$, $e_{\psi,j}=0$, $v_{j}=v_{\text{ref}}$. Formulation (5) bears the advantage of eliminating state dependency and is therefore our preferred form for spatial rate constraints. In practice, we employ bounds that are constant over the spatial planning horizon, but in general time-varying in any closed-loop receding horizon control-setting and dependent on the vehicle’s operating point.
To summarise this section, the transformation of time-dependent control rate constraints (3) to the path-aligned coordinate frame is not trivial and to be considered as the main disadvantage of a spatial-based system representation. We opted for the simple state-independent form (5) for linearly constrained optimisation problems, and discussed the role of $v_{\text{ref}}>0$ as a transformation parameter.
@@ 3.2 Main Two-Step Hierarchical Algorithm
An edgy path segment can either be a headland path segment or a headland-to-mainfield lane transition. In all of the following the term edgy path segment shall refer to path segments with sharp corners or abrupt directional changes. Edgy path segments are identified by heuristics, in particular, by searching for abrupt directional changes above a certain threshold as well as the labeling of path coordinates as headland path or mainfield lane and detecting transitions between the two with sharp directional changes.
@@ 3.3 Problem 1: Headland path smoothing
In Fig. 2 an illustrative edgy headland path segment is labeled with Problem Type 1. The two steps of the hierarchical algorithm for Problem 1 are visualised in Fig. 4: (a) reference path generation and (b) linear programming solution.
In a first step, the segment is scaled by heuristics and replaced by a 5-point piecewise-affine (PWA) path as shown in Fig. 4(a). Several comments are made. First, the key characteristic of this scaling is that the tip of the resulting 5-point PWA path shall ensure zero spraying gap. Thus, the distance between tip and its projection point along the field contour must be less or equal than half the machinery operating width $w>0$. Second, beyond this key characteristic the shaping of the reference path is a heuristic and free choice. Here, a 5-point PWA path segment is used because it rounds out the original edgy headland path segment (more than e.g. an alternative 3-point PWA path segment would) and thereby helps the downstream linear program formulation by providing an improved reference around which system dynamics are linerised and discretised. Third, the heuristic nature of this first reference path generation step is acknowledged and mentioned in the Conclusion as a potential avenue of future research. Fourth and as a detail, we do not use popular Chaikin’s algorithm (see Chaikin (1974)) for corner cutting and further smoothing of the reference path, since this would replace every edge point with two edge points at each iteration step. Furthermore, it would cut off a significant portion of the tip of the 5-point PWA line after the first iteration.
In a second step, using the 5-point PWA path from Step 1 as reference path the following basic LP is proposed:
@@
Steering actuation constraints are (6c) and (6d) as discussed in Sect. 3.1. In practice, constraints (6b) which enforce the resulting path to stay on one side of the reference path are relaxed to, $e_{y,j}-e_{y,j}^{\text{ref}}\leq\sigma$, by introducing a non-negative slack variable $\sigma\geq 0$ and adding it to the objective function 6a with a high weight such that $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|$ is replaced with $\sum\nolimits_{j=1}^{N}|e_{y,j}-e_{y,j}^{\text{ref}}|+10^{16}\sigma$. Furthermore, for the absolute value arguments $N$ non-negative surrogate variables are introduced and the set of inequality constraints is extended accordingly. In the end, there are $n_{u}=2N+1$ scalar real-valued optimisation variables. Typical sizes of the LP (i.e., the number of variables and inequality constraints) are stated in the numerical example Section 4.
@@
Finally, LP-iterations or refinement steps can be conducted. Thus, for the first solution of (6) a 5-point-PWA path is used as reference. Then, the resulting path is used as reference for the second solution of (6). In theory, many of such LP-iterations can be carried out. In practice, one refinement step was found to be useful and sufficient.
@@ 3.4 Problem 2: Headland-to-mainfield lanes transitions
Fig. 6 is of special interest for two reasons. First, it illustrates how Dubins paths are not suited for the generation of precision paths to be followed by autonomous machines online, especially when calculating Dubins paths with minimal turning radius of the machine (see Fig. 6(a)). Due to actuation constraints deviations would be expected in any online tracking application. Second, it illustrates the effect on tracking performance when using different turning radii $R_{\text{Dubins}}$ in the Dubins reference path generation step. When using the minimal turning radius Fig. 6(a) results. By using an artificially larger $R_{\text{Dubins}}$ the LP-solution trajectory follows more closely the reference path, see Fig. 6(b) and 6(c). The drawback of a larger $R_{\text{Dubins}}$ is that the resulting path is longer and more compacted area occurs from corresponding tyre traces and the corresponding steering maneuvers, which reduces growable crop area in the field area. A trade-off with a preference towards small $R_{\text{Dubins}}$ for the reference path generation step is recommended.
@@ 4 Results
Fig. 8: Experiment: 3 instances of Problem 1 and 16 instances of Problem 2 needed to be solved. The path obtained from proposed method is visualised. The sequence of nodes, $[0,16,\dots]$, gives rise to a full field coverage path plan and is calculated by the routing algorithm from Plessen (2019b) subject to the constraint that a full initial headland coverage (from node 00 to node 00 along the headland path before servicing the mainfield lanes) is sought as part of the full area coverage path. The field entrance is denoted by node 00. The inter-lane distance is 20m.
@@
For numerical evaluation the field from Fig. 1 is considered. It has a size of 4.5ha and requires the solution of 3 instances of Problem 1 and 16 instances of Problem 2. The field under consideration involves freeform mainfield lanes, and multiple tight edges for transitions to and from the mainfield lanes and along the headland path.
@@
Fig. 9 shows a zoom-in for an alternative solution approach to Problem 1. Here, a Bezier-curve (Pastva (1998)) of third order is fitted to the 5-point PWA path. A benefit is its simplicity, requiring only the solution of a least-squares problem to identify parameters before the evaluation of a Bernstein polynomial. A disadvantage is that eventhough the interpolating fit is smooth this method does not take vehicle actuation limits into account. The path is therefore in general not feasible to drive. In contrast, proposed LPs (6) and (7) do account for vehicle actuation limits.
@@
Refer to caption Refer to caption
Refer to caption Refer to caption
Fig. 12: More illustrative examples with edgy field contours: 3 fields with straight mainfield lanes and 1 field with freeform curved mainfield lanes. For each example, multiple instances of Problem 1 and Problem 2 had to be solved. Overall, 16 for Problem 1 and 68 for Problem 2.
More illustrative examples are provided in Fig. 12.
5 Discussion
A general comment about importance of input data quality is made. The quality of results for data-dependent techniques is a function of input data. For very edgy field contour input data, as is the case for all 5 fields in Fig. 8 and 12, a decision has to be made: (i) achieving full area coverage but exceeding contours when planning a nonholonomic path (see Fig. 9), or (ii) permitting area coverage gaps but achieving a path with sprayed area fully inside field contours. Both objectives cannot be achieved simultaneously. According to the formulations of Problem 1 and 2 avoiding area coverage gaps is the priority of this paper. Objective (ii) is likewise not trivial and subject of future work.
By counter-examples and discussion, below section is meant to provide further supporting arguments for proposed method.